Diffusion Morning Report 02/13/06

Ted Berg tberg at snf.stanford.edu
Mon Feb 13 06:50:30 PST 2006


Good Morning All,
    Looking fair this morning. Quite a few problems over the weekend. 
Epis pump was off -low oil, TEOS ampule needs changing,
Torch on Tylan 3 replaced-bad uniformity, Tylan 1 had TCU error-testing, 
Tylan2 had MFS error-testing. Everything else looks OK for now.

  	Date: 	2/11/2006 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	yellow 	Had TCU error 	 
Tylan 2 	yellow 	Had MFS alarm 	 
Tylan 3 	yellow 	bad uniformity-changing torch
	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	Down 	teos change 	 
Tylan 8-SiGe 	Up 	  	16.46/30
Tylan 9-Poly 	Up 	  	16.67/30
Tylan 10-Nitride 	Up 	  	2.17/30
Tylan 11-BPSG-LTO 	up 	  	12.81/30
Tylan 12-FGA 	Up 	  	 
  	  	  	 
Thermco 1 	Down 	torch problem? 	 
Thermco 2 	Up 	  	 
Thermco Poly 	Up 	  	 
Thermco Nitride 	Down 	Testing 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	  
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	up 	  	 
  	  	  	 
SiH4 for furnaces 	  	A on line 250 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060213/3e651807/attachment.html>


More information about the diffusion mailing list